OpenStudy (anonymous):

can you help me derive this equation to a linear equation? 1/f = 1/s+1/s'

OpenStudy (anonymous):

\[\frac{ 1 }{ f } = \frac{ 1 }{ s }+\frac{ 1 }{ s' }\]

OpenStudy (anonymous):

from my table, i got this graph of m against s'